OpenStudy (anonymous):

You invest an initial $1,000 in an account that has an annual interest rate of 2%, compounded quarterly. How much money will you have in the account after 5 years? Round your answer to the nearest whole number.

OpenStudy (anonymous):

help me please!

OpenStudy (kropot72):

\[\large A=P(1+\frac{r}{n})^{nt}\] where A is the amount of the principal P after t years, r is the annual interest rate as a decimal fraction and n is the number of compounding periods per year. Do you follow so far?

OpenStudy (anonymous):

yes, i'm following you.

OpenStudy (kropot72):

Good. Can you now plug the given values into the equation and post the result?

OpenStudy (anonymous):

i'm having some trouble doing this...

OpenStudy (kropot72):

What part are you having trouble with?

OpenStudy (anonymous):

plugging the numbers and such in...

OpenStudy (kropot72):

Well I'll plug them in for you, and hopefully you'll see that its not very difficult. \[\large A=1000(1+\frac{0.02}{4})^{(4\times5)}\]

OpenStudy (anonymous):

thank you!

OpenStudy (kropot72):

You're welcome :)

OpenStudy (anonymous):

:)